OpenStudy (anonymous):

You are making poster board boundary markers in the shape of square pyramids for a relay race. If you are including an 8-inch square as the bottom and you want the slant height to be twenty-nine inches, how much poster board do you need? five hundred twenty-eight square inches five hundred seventy-six square inches six hundred forty square inches seven hundred sixty-eight square inches

OpenStudy (anonymous):

A = 1/2bh A=1/2(8)(29) A = 116 (116)(4 triangles) = 464 SA = 464 + 64 = 528 = five hundred twenty - eight square inches.

OpenStudy (anonymous):

Thank you.......

OpenStudy (anonymous):

Yw